I am in pre-algebra. My teacher gave me a worksheet with word problems, but he did not explain how to do it.Can you please help me to understand this? Thank you!
Simplifying algebraic expressions...
This nut bread contains: 2 walnuts, twice as many pecans as walnuts, 10 nuts in all.
I am suppose to find out how many of each type of nut does the bread have.

Answer by stanbon(75887)

This nut bread contains: 2 walnuts, twice as many pecans as walnuts, 10 nuts in all.
I am supposed to find out how many of each type of nut does the bread have.
---------------------
walnuts: 2
pecans: 2*2 = 4
other nuts: 10-(2+4) = 10-6 = 4
===================================
Cheers,
Stan H.
=============